hartley233:

Abbi plans to spend $24 for cashews and peanuts for her party. The Nutty Shop sells peanuts for $3 per pound while cashews are $5 per pound. How many pounds of peanuts can she get if she gets 5 pound total.

ScamTheMan:

Well, this is actually a simple problem using system of equations. You can set up your equations like this: C = Cashews P = Peanuts Since ever peanut is $3 and every cashew is $5 and their total is $24, we can use the expression 3p + 5c = 24. Since both p and c have to equal a pound, we can also use the equation p + c = 5. Our system is now: 3p + 5c = 24 p + c = 5 Do you know the next step from here?